A 5.90-g lead bullet traveling at 510 m/s is stopped by a large tree. If half the kinetic energy of the bullet is transformed into internal energy and remains with the bullet while the other half is transmitted to the tree, what is the increase in temperature of the bullet?
___________
Answer #1

here , specific heat of lead ,

Slead = 0.128 J/gm*K

Now,

as half kinetic energy is converted to heat of bullet

0.5 * 0.5 * m * 510^2 = m* 0.128 * del T

del T = 5.08 *10^5 K

del T = 5.08 *10^50 C

the increase in temperature of the bullet is 5.08 *10^50 C